On September 1, 2019, Jay Walker Company purchased a one-year insurance policy for $1,320. The correct adjusting entry on December 31, 2019, is:

Multiple Choice

  • debit Insurance Expense $440; credit Prepaid Insurance $440

  • debit Insurance Expense $990; credit Prepaid Insurance $990

  • debit Prepaid Insurance $110; credit Insurance Expense $110

  • debit Prepaid Insurance $1,320; credit Insurance Expense $1,320
Answer #1

Answer is debit Insurance Expense $440; credit Prepaid Insurance $440

Insurance expense =$1320/12 months * 4 months =$440
Answer #2

Calculation:

  1. Total Insurance Cost: $1,320 for 12 months.

  2. Monthly Insurance Expense: 1,320÷12=110 per month .

  3. Covered Period (Sep 1 - Dec 31): 4 months.

  4. Expense for 4 Months: 4 × 110=440 .


Adjusting Entry:

  • Debit Insurance Expense $440 (recognize the expense for 4 months).

  • Credit Prepaid Insurance $440 (reduce the prepaid asset by the used amount).



Answer:

  • debit Insurance Expense 440;creditPrepaidInsurance

The remaining prepaid insurance after adjustment: 1,320440 = $880 (for the next 8 months)


answered by: Harshwardhan kunal